The length of the rectangle is 2m more than the width. The area of the rectangle is 48m squared. Find the length and the width.
Answer by abigal95(2)

the equation to start with is
L x W = A
you are given that w+2=L
so substitute w+2 for L in the original equation and plug in the value of the area given
w(w+2)=48
and foil it out and subtract 48 from both sides to get
w^2+2w-48=0
then you will have to factor the equation which comes out to be
(w+8)(w-6)
the (W+8) will be set equal to 0 and w will be negative so then (W-6) will be set equal to 0 and then that equals 6 which is the width and then the length would be 8 since W+2 equals the length, and 6 times 8 is 48